OpenStudy (anonymous):

what is 3y-x=-6 in slope intercept form?

OpenStudy (anonymous):

-y/2 + x/6 = 1

OpenStudy (anonymous):

divide whole thing by what given on right hand of equality

OpenStudy (anonymous):

okayy so do you know what form the "slope intercept" is in?? @brandon10995

OpenStudy (anonymous):

y=mx+b yes i know that. Just confused with this one

OpenStudy (anonymous):

oops sorry .... that will be y=x/3 -2

OpenStudy (anonymous):

:)) okay 3y-x=-6. we have to isolate y. we could divide everything by 3 to get the y alone

OpenStudy (anonymous):